OpenStudy (anonymous):

find the missing exponent. 2^ =1/4

OpenStudy (anonymous):

.8

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \sf 2^x=1/4 }\) \[\Huge\color{blue}{ \sf \log(2^x)=\log(1/4) }\] \[\Huge\color{blue}{ \sf xlog(2)=\log(1/4) }\] \[\Huge\color{blue}{ \sf x=\log(1/4) \div \log(2) }\] \[\Huge\color{blue}{ \sf x=\log_2(1/4) }\]

OpenStudy (solomonzelman):

I don't have a calculator with me, sorry.